OpenStudy (anonymous):

sketch 2 periods of the following function: y=tan(2x)+3. showing all critical points. I mostly need help with labeling the x-axis. I always get confused when there's a phase shift.

OpenStudy (anonymous):

critical points? you doing caculus now?

OpenStudy (anonymous):

yea i have a final exam tomorrow and this is one of the things i don't understand

OpenStudy (anonymous):

basically she wants us to show all the points for the whole 2 periods

OpenStudy (anonymous):

there is no phase shift here because it is \(\tan(2x)\) and not \(\tan(2x+c)\)

OpenStudy (anonymous):

tangent is periodic with period \(\pi\) and has asymptotes at \(y=-\frac{\pi}{2}\) and \(y=\frac{\pi}{2}\)

OpenStudy (anonymous):

if you want your new asymptotes, rather than memorizing some formula, set \(2x=-\frac{\pi}{2}\) and \(2x=\frac{\pi}{2}\) and solve for \(x\). that will give them to you

OpenStudy (anonymous):

you will see that your function looks just like tangent, only instead of having period \(\pi\) it has period \(\frac{\pi}{2}\) so you might want to label the \(x\) axis at the points \(-\frac{\pi}{4},\frac{\pi}{4}, \frac{\pi}{2}\) and graph the two periods there

OpenStudy (anonymous):

of course it is also shifted up 3 units because of the \(+3\) at the end here is a nice picture, notice the units along the x axis http://www.wolframalpha.com/input/?i=tan%282x%29%2B3

OpenStudy (anonymous):

why did u set up \[\Pi/2=2x\]

OpenStudy (anonymous):

@satellite73 you there?

OpenStudy (anonymous):

i set it up because i know the basic graph of \(y=\tan(x)\) it has asymptotes at \(x=-\frac{\pi}{2}\) and \(x=\frac{\pi}{2}\)

OpenStudy (anonymous):

here is what i have in my mind http://www.wolframalpha.com/input/?i=tan%28x%29

OpenStudy (anonymous):

so starting with that basic graph, we now translate with \[y=\tan(2x)+3\] the \(+3\) at the end means we shift up 3 units as for the asymptotes, we have \(2x\) not \(x\) as the input, so i set \(2x=-\frac{\pi}{2}\) solve for \(x\) get \(x=-\frac{\pi}{4}\) so that is the left hand asymptote

OpenStudy (anonymous):

right hand endpoint i put \(2x=\frac{\pi}{2}\) get \(x=\frac{\pi}{4}\) so that is the right hand asymptote this also tells us the period is \(\frac{\pi}{2}\) the distance between the two points. we could halve also used that the period of \[\tan(bx)\] is \(\frac{\pi}{b}\)

OpenStudy (anonymous):

*have

OpenStudy (anonymous):

oh ok, i think i get it

OpenStudy (anonymous):

check out both links i sent and make sure it is clear how to get from the second one, the basic graph, to the first one, your answer

OpenStudy (anonymous):

ok, i mostly don't understand why the graph changed from \[\Pi/2\] to \[\Pi/4\]

OpenStudy (anonymous):

ok you do understand that one asymptote of tangent is at \(\frac{\pi}{2}\) right? because tangent is sine over cosine, and cosine is undefined there

OpenStudy (anonymous):

but you do not have \(\tan(x)\) you have \(\tan(2x)\)

OpenStudy (anonymous):

so suppose \(x=\frac{\pi}{4}\) then \[\tan(2x)=\tan(2\times\frac{\pi}{4})=\tan(\frac{\pi}{2})\] is undefined. so that is your new asymptote

OpenStudy (anonymous):

why is it undefined? i know tangent is sine over cosine but i'm still confused

OpenStudy (anonymous):

another way to think about it is that as \(x\) goes from \(-\frac{\pi}{4}\) to \(\frac{\pi}{4}\) we have \(2x\) goes from \(-\frac{\pi}{2}\) to \(\frac{\pi}{2}\) one full period for tangent

OpenStudy (anonymous):

ok \[\tan(x)=\frac{\sin(x)}{\cos(x)}\] \[\tan(\frac{\pi}{2})=\frac{\sin(\frac{\pi}{2})}{\cos(\frac{\pi}{2})}=\frac{1}{0}\]is undefined

OpenStudy (anonymous):

because you cannot divide by 0 so \(\frac{\pi}{2}\) is not in the domain of tangent, and that is why it has a vertical asymptote there

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

tangent is undefined anywhere cosine is 0 so it is undefined at \(-\frac{\pi}{2}, \frac{\pi}{2}, \frac{3\pi}{2}, ...\)